OpenStudy (anonymous):

wats 2x+4=x+6 and 2x=x+2. please show work and explane.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

2x=x+2 -x -x x=2

OpenStudy (anonymous):

2x+4=x+6 x+4=6 x=2